Vapour density of metal chloride is 66. Its oxide contains 53% metal . What is the atomic mass of the metal ?

Grade:

1 Answers

Arun
5 years ago
2×V.D= molecular mass.
.:2×66=132 is moecular mass of the metal.
its oxide has 53%metal. i.e 100 gm oxide will have 53 gm metal.that is 47 gm of Oxygen i.e. approximatly 3 moles of oxygen. 
So one molecule of metal has 3 atoms of oxygen. This implies that it is of form
A2O3. 
Most probably Al2O3.
Now let's cross chek.
atomic mass of Al is 26.5 
26.5×2= 53.
(Yeah we got it!)
Its chloride is AlCl3.
26.5+3×35.5= 26.5+ 106.5=133.
So hete is your answer. atomic wt. of the metal is 26.5.